# Improving LLM Unlearning Robustness via Random Perturbations

### v3

@@ 3 A Unified View of LLM Unlearning
$\mathbb{E}_{\mathbf{v}}[\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v})]$ $\approx\mathbb{E}_{\mathbf{v}}[\ell(y^{f}|\mathbf{z}_{\theta}^{f})]+\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})]+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})\mathbf{v}]$ (6)
$=\ell(y^{f}|\mathbf{z}_{\theta}^{f})+\nabla_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})^{\top}\mathbb{E}_{\mathbf{v}}[\mathbf{v}]+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})\mathbf{v}]$ (7)
$=\ell(y^{f}|\mathbf{z}_{\theta}^{f})+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})\mathbf{v}],\quad\text{since}\;\mathbb{E}_{\mathbf{v}}[\mathbf{v}]=\bm{0}.$ (8)
A classic result from Hutchinson [24] tell us that $\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})\mathbf{v}]=\mu\text{Tr}(\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f}))$, where $\text{Tr}(\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f}))>0$ is the trace of the positive definite Hessian matrix $\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})$ (by assumption). Since $\mu\in\mathbb{R}^{+}$, the loss of generating $y^{f}$ given latent variable $\mathbf{z}_{\theta}^{f}$ is increases, that is,
$\mathbb{E}_{\mathbf{v}}[\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v})]$ $\approx\ell(y^{f}|\mathbf{z}_{\theta}^{f})+\frac{\mu}{2}\text{Tr}(\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f}))>\ell(y^{f}|\mathbf{z}_{\theta}^{f})$ (9)
While presenting in different formulations, PO and RM share a common high-level principle—maximizing the loss of forget-samples $\ell(y^{f}|\mathbf{z}_{\theta}^{f})$. Therefore, Eqn. 9 suggests that steering forget-representations toward a random representation in RM is effectively equivalent to maximizing the loss of those forget-samples in PO. In other words, PO can be viewed as RM, that is, PO introduces noise-like effects to the forget-representation during fine-tuning, disrupting its alignment with target labels. We present an empirical validation in Appendix C.
@@ 4.1 Threat Model
We define the threat model and the unlearning guarantee that is expected to hold. We consider a practical scenario, such as machine learning as a service (MLaaS), where users can black-box access the unlearned model through an API.
@@
Unlearning guarantee. The unlearned models are expected to be robust against forget-tokens in retain-queries while maintaining the forgetting performance on forget-tasks as well as retaining performance on benign retain-queries. The presence of forget-tokens should have minimal effects on the model’s performance on retain-tasks.
@@ 4.2 Robustness of Unlearned Models Against Forget-Tokens
Let $x^{r}_{i}$ be the generated token given the previous retain sequence $x^{r}_{<i}$ in the unlearned model $f^{u}$. $x^{r,\text{per}}_{<i}$ denotes the perturbed retain-query (the retain-query that contains forget-tokens). Let $\mathbf{z}^{r}_{<i}$ and $\mathbf{z}^{r,\text{per}}_{<i}$ be the representation of $x^{r}_{<i}$ and $x^{r,\text{per}}_{<i}$ respectively obtained from $f^{u}$ at layer $l$. To formalize the analysis, we introduce the following assumption.
@@
Assumption 4.1 implies that the presence of forget-tokens in the retain-query introduces uncertainty in the model’s latent representations. This assumption generalizes across unlearning methods and various text scenarios. The scalar $\eta$ controls the magnitude of perturbations capturing the variation of forget-tokens can be appeared in the perturbed retain-queries. Based on this assumption, we derive the change in the output representation of the generated tokens as follows.
@@
If Assumption 4.1 holds, the change in the output representation of the generated token $x^{r}_{i}$ given the perturbed retain-query $x^{r,\textnormal{per}}_{<i}$ and the benign retain-query $x^{r}_{<i}$ in the unlearned model $f^{u}$, defined as $\Delta=f^{u}(x^{r}_{i}|x^{r,\text{per}}_{<i})-f^{u}(x^{r}_{i}|x^{r}_{<i})$, follows the Normal distribution $\mathcal{N}(\bm{0},\eta\bm{J}^{\top}\bm{J})$, where $\bm{J}=\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(x^{r}_{i}|x^{r}_{<i})$ is the Jacobian of $f^{u}(x^{r}_{i}|x^{r}_{<i})$ with respect to $\mathbf{z}^{r}_{<i}$.
@@
5 Random Noise Augmentation
5.1 Motivation: Unlearning as a Backdoor Attack and Defense Problem
As a motivation, our first start is an assumption that the unlearning process can be decomposed into two parts: the “forgetting” process and the “retaining” process. We then reframe the unlearning as a backdoor attack and defense problem and propose a simple yet effective solution to enhance the robustness of unlearned models.
@@ 5.1 Motivation: Unlearning as a Backdoor Attack and Defense Problem
“Forgetting” as a backdoor attack. Consider the supervised learning setting with the objective of learning a model $f_{\theta}:\mathcal{X}\mapsto\mathcal{Y}$. Let $\mathcal{Z}=\mathcal{Z}_{f}\cup\mathcal{Z}_{r}$ be the “latent representation” dataset corresponding to the original dataset $\mathcal{D}=\mathcal{D}_{f}\cup\mathcal{D}_{r}$. $\mathcal{Z}$ is composed of a forget-set $\mathcal{Z}_{f}=\{(\mathbf{z}_{\theta}^{f},\mathbf{z}_{\theta^{\text{ref}}}^{f})\}_{i}$, where $\mathbf{z}_{\theta}^{f}\in\mathcal{X}$ is the input, $\mathbf{z}_{\theta^{\text{ref}}}^{f}\in\mathcal{Y}$ is the target output, and a retain-set $\mathcal{Z}_{r}=\{(\mathbf{z}_{\theta}^{r},\mathbf{z}_{\theta^{\text{ref}}}^{r})\}_{j}$ where $\mathbf{z}_{\theta}^{r}\in\mathcal{X}$ and $\mathbf{z}_{\theta^{\text{ref}}}^{r}\in\mathcal{Y}$. Each forget-sample $(\mathbf{z}_{\theta}^{f},\mathbf{z}_{\theta^{\text{ref}}}^{f})$ is transformed into a backdoor-sample $(T(\mathbf{z}_{\theta}^{f}),\Omega(\mathbf{z}_{\theta^{\text{ref}}}^{f}))$, where $\Omega$ is an adversarial-target labeling function and $T$ is the trigger generation function. In a standard backdoor attack, $T$ is usually optimized for generating and placing the trigger into the input while $\Omega$ specifies the behavior of the model when the backdoor trigger is activated. In the “forgetting”, $T$ is an identity function i.e. $T(\mathbf{z}_{\theta}^{f})=\mathbf{z}_{\theta}^{f}$ and $\Omega$ is a function that maps $\mathbf{z}_{\theta^{\text{ref}}}^{f}$ to the adversarial-perturbed representation (e.g., $c\mathbf{u}$ in RMU). We train model $f_{\theta}$ with “poisoned” forget-set $\mathcal{Z}^{\text{poisoned}}_{f}=\{(T(\mathbf{z}_{\theta}^{f})),\Omega(\mathbf{z}_{\theta^{\text{ref}}}^{f}))\}_{i}$ and benign retain-set $\mathcal{Z}_{r}=\{(\mathbf{z}_{\theta}^{r},\mathbf{z}_{\theta^{\text{ref}}}^{r})\}_{j}$, as follows:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim\mathcal{Z}}\left[\mathcal{L}(f_{\theta}(\mathbf{x}),\mathbf{y})\right],$ (10)
@@
where $\mathbf{x}$ is either $\mathbf{z}_{\theta}^{f}$ or $\mathbf{z}_{\theta}^{r}$ and $\mathbf{y}$ is either $\mathbf{z}_{\theta^{\text{ref}}}^{f}$ or $\mathbf{z}_{\theta^{\text{ref}}}^{r}$. During inference, for a retain-input $\mathbf{z}_{\theta}^{r}$ and forget-input $\mathbf{z}_{\theta}^{f}$ the unlearned model should behave as follows:
$f(\mathbf{z}_{\theta}^{r})$ $=\mathbf{z}_{\theta^{\text{ref}}}^{r}$ (11)
$f(\mathbf{z}_{\theta}^{f})$ $=f(T(\mathbf{z}_{\theta}^{f}))=\Omega(\mathbf{z}_{\theta^{\text{ref}}}^{f})$ (12)
@@
This formulation suggests that current state-of-the-art LLM unlearning methods themselves “poison” the model and make it more vulnerable to forget-tokens. The presence of the forget-token in the retain-queries is equivalent to the activation of the backdoor trigger in these queries, leading the model to misbehave. This backdoor explanation further highlights the evidence that current LLM unlearning methods do not truly erase knowledge; in fact, they intentionally decide that the model’s target knowledge/behaviors should not be surfaced [29, 10].
@@
“Retaining” as a backdoor defense. We then came up with an idea to treat the “retaining” as a backdoor defense. The goal is to reduce the sensitivity of the unlearned models to forget-tokens. Inspired by previous works using random noise defenses against adversarial attacks [31, 19, 48, 45, 6, 23], we propose Random Noise Augmentation (RNA), adds a small, independent random Gaussian noise $\bm{\delta}\sim\mathcal{N}(\bm{0},\nu\bm{I})$, $\nu\in\mathbb{R}^{+}$ to each retain-representation in the reference model during fine-tuning.
@@
Algorithm. The process of RNA is described in Algorithm 1. The core intuition behind incorporating randomness into the latent space of the model aims to confuse the “backdoor attacker” and steer it away from its “unintended” objectives on retain-queries. Notably, RNA offers several compelling advantages: (1) RNA is plug-and-play, model- and method-agnostic: RNA can be applied to any deep networks and generalizes to the commonly used form of MU, especially to the two unlearning frameworks, including RM and PO. After the forward pass, the randomized logit and representation of the retain-sample in the reference model can be used as the target retain output in the retain-loss of PO and RM, respectively.
@@
0:  a $L$-layer reference model $f_{\theta^{\text{ref}}}$, a retain-sample $x^{r}$, a layer $l\in[1...L]$, a noise scale $\nu$.
0:  return logit and representation of $x^{r}$.
@@
4:        $\mathbf{z}^{r}_{\theta^{\text{ref}}}\leftarrow\mathbf{z}^{r}_{\theta^{\text{ref}}}+\bm{\delta}$.

### v4

@@ 3 A Unified View of LLM Unlearning
$\mathbb{E}_{\mathbf{v}}[\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v})]$ $\approx\mathbb{E}_{\mathbf{v}}[\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})]+\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})]+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})\mathbf{v}]$ (6)
$=\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})+\nabla_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\theta}^{f})^{\top}\mathbb{E}_{\mathbf{v}}[\mathbf{v}]+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})\mathbf{v}]$ (7)
$=\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})+\frac{1}{2}\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})\mathbf{v}],\quad\text{since}\;\mathbb{E}_{\mathbf{v}}[\mathbf{v}]=\bm{0}.$ (8)
A classic result from Hutchinson (1989) tell us that $\mathbb{E}_{\mathbf{v}}[\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})\mathbf{v}]=\mu\text{Tr}(\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}))$, where $\text{Tr}(\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}))>0$ is the trace of the positive definite Hessian matrix $\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})$ (by assumption). Since $\mu\in\mathbb{R}^{+}$, the loss of generating $\mathbf{y}^{f}$ given latent variable $\mathbf{z}_{\bm{\theta}}^{f}$ is increases, that is,
$\mathbb{E}_{\mathbf{v}}[\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v})]$ $\approx\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})+\frac{\mu}{2}\text{Tr}(\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}))>\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})$ (9)
While presenting in different formulations, PO and RM share a common high-level principle—maximizing the loss of forget-samples. Therefore, Eqn. 9 suggests that steering forget-representations toward a random representation in RM is effectively equivalent to maximizing the loss of those forget-samples in PO. In other words, PO can be viewed as RM, that is, PO introduces noise-like effects to the forget-representation during fine-tuning, disrupting its alignment with target labels. We present an empirical validation in Appendix C.1.
@@ 4.1 Threat Model
We first define the threat model and the unlearning guarantee that is expected to hold. We consider a practical scenario, such as machine learning as a service (MLaaS), where users can black-box access the unlearned model through an API.
@@
Unlearning guarantee. Unlearned models are expected to be robust against forget-tokens in retain-queries while maintaining the forgetting performance on forget-tasks as well as retaining performance on benign retain-queries. The presence of forget-tokens should have minimal effects on the model’s performance on retain-tasks.
@@ 4.2 Robustness of Unlearned Models Against Forget-Tokens
Let $\mathbf{x}^{r}_{i}$ be the generated token given the previous retain sequence $\mathbf{x}^{r}_{<i}$ in the unlearned model $f^{u}$. $\mathbf{x}^{r,\text{per}}_{<i}$ denotes the perturbed retain-query (the retain-query that contains forget-tokens). Let $\mathbf{z}^{r}_{<i}$ and $\mathbf{z}^{r,\text{per}}_{<i}$ be the representation of $\mathbf{x}^{r}_{<i}$ and $\mathbf{x}^{r,\text{per}}_{<i}$ respectively obtained from $f^{u}$ at layer $l$. To formalize the analysis, we make the following assumption.
@@
Assumption 4.1 implies that the presence of forget-tokens in the retain-query introduces uncertainty in the model’s latent representations. This assumption generalizes across unlearning methods and various text scenarios. The scalar $\eta$ controls the magnitude of perturbations, capturing the variation of forget-tokens that can appear in the perturbed retain-queries. Next, we derive the change in the output representation of the generated tokens as follows.
@@
If Assumption 4.1 holds, the change in the output representation of the generated token $\mathbf{x}^{r}_{i}$ given the perturbed retain-query $\mathbf{x}^{r,\text{per}}_{<i}$ and the benign retain-query $\mathbf{x}^{r}_{<i}$ in the unlearned model $f^{u}$, defined as $\Delta=f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r,\text{per}}_{<i})-f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r}_{<i})$, follows the Normal distribution $\mathcal{N}(\bm{0},\eta\bm{J}^{\top}\bm{J})$, where $\bm{J}=\nabla_{\mathbf{z}^{r}_{<i}}f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r}_{<i})$ is the Jacobian of $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r}_{<i})$ with respect to $\mathbf{z}^{r}_{<i}$.
@@
5 Unlearning as A Backdoor Attack and Defense Problem
5.1 Formulation
“Forgetting” as a backdoor attack.
@@ 5.1 Formulation
We formulate the “Forgetting” process as a backdoor attack. Consider the supervised learning setting with the objective of learning a model $f_{\bm{\theta}}:\mathcal{X}\mapsto\mathcal{Y}$. Let $\mathcal{Z}=\mathcal{Z}_{f}\cup\mathcal{Z}_{r}$ be the “latent representation” dataset corresponding to the original dataset $\mathcal{D}=\mathcal{D}_{f}\cup\mathcal{D}_{r}$. $\mathcal{Z}$ is composed of a forget-set $\mathcal{Z}_{f}=\{(\mathbf{z}_{\bm{\theta}}^{f},\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f})\}_{i}$, where $\mathbf{z}_{\bm{\theta}}^{f}\in\mathcal{X}$ is the input, $\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f}\in\mathcal{Y}$ is the target output, and a retain-set $\mathcal{Z}_{r}=\{(\mathbf{z}_{\bm{\theta}}^{r},\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{r})\}_{j}$ where $\mathbf{z}_{\theta}^{r}\in\mathcal{X}$ and $\mathbf{z}_{\theta^{\text{ref}}}^{r}\in\mathcal{Y}$. Each forget-sample $(\mathbf{z}_{\bm{\theta}}^{f},\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f})$ is transformed into a backdoor-sample $(T(\mathbf{z}_{\theta}^{f}),\Omega(\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f}))$, where $\Omega$ is an adversarial-target labeling function and $T$ is the trigger generation function. In a standard backdoor attack, $T$ is usually optimized for generating and placing the trigger into the input while $\Omega$ specifies the behavior of the model when the backdoor trigger is activated. In the “forgetting”, $T$ is an identity function i.e. $T(\mathbf{z}_{\bm{\theta}}^{f})=\mathbf{z}_{\bm{\theta}}^{f}$ and $\Omega$ is a function that maps $\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f}$ to the adversarial-perturbed representation (e.g., $c\mathbf{u}$ in RMU). We train model $f_{\bm{\theta}}$ with “poisoned” forget-set $\mathcal{Z}^{\text{poisoned}}_{f}=\{(T(\mathbf{z}_{\bm{\theta}}^{f})),\Omega(\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f}))\}_{i}$ and benign retain-set $\mathcal{Z}_{r}=\{(\mathbf{z}_{\bm{\theta}}^{r},\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{r})\}_{j}$, as follows:
$\bm{\theta}^{*}=\arg\min_{\bm{\theta}}\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim\mathcal{Z}}\left[\mathcal{L}(f_{\bm{\theta}}(\mathbf{x}),\mathbf{y})\right],$ (10)
@@
where $\mathbf{x}$ is either $\mathbf{z}_{\bm{\theta}}^{f}$ or $\mathbf{z}_{\bm{\theta}}^{r}$ and $\mathbf{y}$ is either $\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f}$ or $\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{r}$. During inference, for a retain-input $\mathbf{z}_{\bm{\theta}}^{r}$ and forget-input $\mathbf{z}_{\bm{\theta}}^{f}$ the unlearned model should behave as follows:
$f(\mathbf{z}_{\bm{\theta}}^{r})$ $=\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{r}$ (11)
$f(\mathbf{z}_{\bm{\theta}}^{f})$ $=f(T(\mathbf{z}_{\bm{\theta}}^{f}))=\Omega(\mathbf{z}_{\bm{\theta}^{\text{ref}}}^{f})$ (12)
@@
This formulation suggests that current state-of-the-art LLM unlearning methods themselves “poison” the model and make it more vulnerable to forget-tokens. The presence of the forget-token in the retain-queries is equivalent to the activation of the backdoor trigger in these queries, leading the model to misbehave. This backdoor explanation further highlights that current LLM unlearning methods do not truly erase knowledge; in fact, they intentionally decide that the model’s target knowledge/behaviors should not be surfaced.
“Retaining” as a backdoor defense.
@@
We then came up with an idea to treat the “Retaining” as a backdoor defense. The goal is to reduce the sensitivity of the unlearned models to noises caused by forget-tokens. Inspired by adversarial training, we propose Random Noise Augmentation (RNA), a robust unlearning method, which adds a small, independent random Gaussian noise $\bm{\delta}\sim\mathcal{N}(\bm{0},\nu\bm{I})$, $\nu\in\mathbb{R}^{+}$ to each retain-representation in the reference model during fine-tuning. RNA forget-loss enforces forgetting on the forget-set while the retain-loss preserves general performance and encourages retain-robustness against random perturbations.
5.2 Random Noise Augmentation
@@
The process of RNA is described in Algorithm 1. The core intuition behind incorporating randomness into the latent space of the model aims to confuse the “backdoor attacker” and steer it away from its “unintended” objectives on retain-queries. Notably, RNA offers several compelling advantages: (1) RNA is lightweight, model- and method-agnostic: RNA can be applied to any deep networks and generalizes to the most commonly used form of MU, especially to the two unlearning frameworks, including RM and PO. After the forward pass, the randomized logit and representation of the retain-sample in the reference model can be used as the target retain output in the retain-loss of PO and RM, respectively.
@@
0: a $L$-layer reference model $f_{\bm{\theta}^{\text{ref}}}$, a retain-sample $\mathbf{x}^{r}$, a layer $l\in[1...L]$, a noise scale $\nu$.
0: return logit and representation of $\mathbf{x}^{r}$.
@@
4:   $\mathbf{z}^{r}_{\bm{\theta}^{\text{ref}}}\leftarrow\mathbf{z}^{r}_{\bm{\theta}^{\text{ref}}}+\bm{\delta}$.